OpenStudy (anonymous):

Factor first, and then find the roots. x2 – 5x – 6 = 0

OpenStudy (anonymous):

\[x^2-5x-6=0\] \[(x-6)(x+1)=0\] \[x-6=0\] \[x=6\] \[x+1=0\] \[x=-1\]